i) Ken Downing has offered his personal guarantee to support a loan to his business. His personal financial statement shows only three liabilities:a home mortgage with a remaining balance of $124,980 and monthly payments of $900/-; an auto loan with a remaining balance of $5,400 and monthly payments of $300, and a personal loan with single payment of $3,200 due in 18 months. What is the total of Ken Downing's current liabilities (liabilities due within one year)?

A) $8,600

B) $14,400

C) $16,533

D) $17,600

ii) Which of the following lists of personal assets is shown in order of liquidity, from most to least liquid?

A) Retirement saving plan, listed corporate stock, note receivable

B) Listed corporate bonds, retirement saving plan, unlisted stock

C) Cash value of life insurance, tax refund, closed-end mutual fund

iii) A business owner makes alimony and child support payments that are not shown as a liability on his personal financial statement. What is the correct treatment of these payments for your analysis?

A) Decrease cash by the annualized amount of the payments.

B) Increase current liabilities by the annualized amount of the payments.

C) Increase contingent liabilities by the total amount of future payments

D) Disregard these payments because they are not material to your analysis.
Answer #1

i) total of Ken's Downing liabilities

payment of house loan in a year =900*12= $10800

payment of auto loan in a year= 300*12 = $3600

total liabilities = $ 14400

personal loan payment will not be consider hare because it is due in 18 months

ii) option C is arranged in order of liquidity from most liquid to least liquid ( liquidity means which is readily available)

iii) option B increase current liabilities by the annualized amount of the payments.